

********************************************************************
UARS WEEKLY STATUS REPORT - 12 Apr 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE entered day only operations on April 3rd due to low power
conditions. We continue to perform default solar observations during
this time.
***********
SUSIM
***********
SUSIM performed its daily mid-resolution and MgII solar scans during
periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
PEM support of the UARS Traceability Mission for the TIMED mission has
been organized with the GVUI team. Predicted TIMED-UARS conjunctions from
the high geographic latitudes were compared with the expected magnetic position
and the PEM climatology. Positional and anticipated precipitation flux
information showed that expected electron precipitation as a function of
magnetic activity had the highest probably of a successful conjunction on
April 19. Continued monitoring has shown changes in conjunction times. For
conjunctions within 300 km, there are now 5 times shown:
Contact # Date Day Time Geo Lat ILAT MLT Hemisphere
(deg) (deg) (hr)
1 April 18 108 21:29 56.47 54.53 21.37 North
2 April 18 108 22:18 -56.91 65.92 10.58 South
3 April 18 108 23:06 56.93 57.59 22.07 North
4 April 23 113 15:44 -56.97 47.79 9.10 South
5 April 23 113 16:32 56.96 53.25 20.12 North
Contacts are progressing to earlier times. Of coarse, what is important is
when PEM passes through the aurora at the same time GUVI is imaging, and this
does not mean that UARS and TIMED are in conjunction.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 30 March 2002 (Orbit 57679, GMT Day 089) through 05
April 2002 (Orbit 57784, GMT Day 095).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 088/1039 57670 Power OFF
HALOE 088/1334 57672 Power ON
HALOE 089/1918 57691 Power OFF (TBAT TEMP)
HALOE 089/2234 57693 Power ON
ACRIM 093/0312 57741 Power OFF
PEM AXIS 093/0312 57741 Power OFF
The beta angle ranged between the angles of 52.42 degrees to 34.23
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 13 April 2002 (DOY 103).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 24.96 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.31 to 1.63 degrees C. Additional
battery performance data for each battery on 30 March 2002 and 05
April 2002 are provided in Appendix D.
The clock error ranged from -15.9 to 11.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 01
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
HALOE = 04
WINDII = 01
ACRIM = 01
PEM AXIS = 01
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 3
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/043 - 2287.498879
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
At the Eclipse Delta ORR on 04 April 2002, the Eclipse Ground System
was officially approved as the flight operations ground system for
UARS. This was necessary to meet the anticipated loss of MSOCC
personnel on 05 April 2002; thereby, reducing the cost for the
extended mission. Eclipse has been very stable and performing well;
in addition, the flight operators are adapting at an overwhelming
rate. There are only a few open issues that are being investigated;
these as well as a few other minor fixes are currently in progress.
Another database delivery will probably be scheduled in a few weeks,
allowing for other discrepancies to be found and then incorporated
into the planned changes. Efforts are currently being directed to the
readiness of Eclipse for the upcoming yaw maneuver on 13 April 2002.
MEETINGS
The FOT attended a UARS Eclipse Delta ORR on Thursday, April 04, 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 13 April 2002 (DOY 103). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 April 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
57670 088/1039 02-106 HALOE Power OFF
57672 088/1334 02-107 HALOE Power ON
57682 089/0536 02-108 OBC/CLK Clk Adj 20.6 to 0.00 msec/day
57691 089/1918 02-109 HALOE Power OFF (TBAT TEMP)
57693 089/2234 02-110 HALOE Power ON
57710 091/0135 02-111 WINDII WIERRFIX FOR E03 ERROR CODE
57723 091/2232 02-112 OBC Clk Adj fr 0.00 to 20.6 msec/day
57741 093/0312 02/113 ACRIM Power OFF
57741 093/0312 02-114 PEM AXIS Power OFF
57741 093/0351 02-115 SOLSTICE to DAY ONLY Ops
57751 093/1829 02-116 OBC Clk Adj fr 20.6 to 0.00 msec/day
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 April 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
WIERRCODE WINDII Software Error Code 57710
______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 April 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
57682 089/0536 Clk Adj 20.6 to 0.00 msec/day
57723 091/2232 Clk Adj fr 0.00 to 20.6 msec/day
57751 093/1829 Clk Adj fr 20.6 to 0.00 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 April 2002
30 March 2002 - GMT Day 089 Beta = 52.42 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 0.00/-16.8 -3.77/-4.38 +0.0/ -0.0 10.56 0.000 0.00
2 +11.2/-44.8 +5.14/+3.82 +25.2/-14.4 24.96 1.020 20.04
3 +16.8/-56.0 +3.82/+2.85 +24.0/-14.4 24.96 1.020 19.69
05 April 2002 - GMT Day 095 Beta = 34.23 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 728.0/728.0 -3.16/-4.38 +0.0/ -0.0 9.60 0.000 0.00
2 +11.2/-16.8 +4.48/+3.82 +26.8/-13.2 26.40 1.040 18.99
3 +11.2/-22.4 +3.82/+2.85 +25.6/-12.8 26.40 1.030 18.52
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 April 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-014 FOT 57704 090/1653 14170 Generic Late Acquisition
02-023 FOT 57767 094/2027 24285 Early LOS on OMNI Pass
02-024 FOT 57783 095/2209 24285 SOLSTICE INOPS in Day
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-011 FOT 57466 074/2033 N/A AP-7 Disc Drive Failure
02-013 FOT 57467 074/2116 N/A AP-1 Hang-up
02-016 FOT 57523 078/1413 N/A AP-1 Halt/Gateway Fail
02-017 FOT 57553 080/1522 N/A Multiple TAC Problems
02-018 FOT 57556 080/1930 N/A AP-1 Crashed
02-019 FOT 57579 082/0822 N/A AP Reboot
02-020 FOT 57591 083/0325 N/A AP 1 Crashed
02-021 FOT 57604 084/0022 N/A AP 1 Crashed
02-014 FOT See List Mar 02 N/A Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode
02-023 FOT 57767 094/2027 24285 Early LOS on OMNI Pass
02-024 FOT 57783 095/2209 24285 SOLSTICE INOPS in Day